Let f:[a, b] → ℝ be a function and let c ∈ (a, b). If f is of bounded variation on [a, b],

prove that f is bounded on [a, b] and v(f,[a, b] = v(f,[a, c] + v(f,[c,d]).


1
Expert's answer
2021-09-12T17:34:42-0400

Definitions. A finite subset "P=\{x_0, x_1, \dots,x_{n_P}\}\subset[a,b]" is a partition of the segment "[a,b]" if and only if "a=x_0<x_1<\dots,x_{n_P-1}<x_{n_P}=b". The set of all partitions of the segment "[a,b]" is denoted by "Part[a,b]". If "f:[a,b]\to \mathbb{R}" is a function, then the variation of "f(x)" on the partition P is the sum "V(f;P)=\sum\limits_{i=1}^{n_P}|f(x_i)-f(x_{i-1})|". A quantity "V_a^b(f):=\sup\{V(f;P): P\in Part[a,b]\}" is called by the total variation of the function "f(x)" on the segment "[a,b]". The function "f(x)" is a function of bounded variation on the segment "[a,b]", if and only if "V_a^b(f)<+\infty".


Theorem 1. If "f(x)" is a function of bounded variation on the segment "[a,b]", then "f(x)" is bounded on "[a,b]".

Proof. Suppose that "f(x)" is unbounded on "[a,b]". Then for all "k\in\mathbb{N}" there exists "t_k\in [a,b]" such that "|f(t_k)|\to+\infty", as "k\to+\infty" . Consider the partition "P_k=\{a,t_k,b\}\in Part[a,b]".

"V(f;P_k)=|f(t_k)-f(a)|+|f(t_k)-f(b)|\geq 2|f(t_k)|-|f(a)|-|f(b)|\to+\infty"

Therefore, "V_a^b(f)=\sup\{V(f;P): P\in Part[a,b]\}=+\infty" and "f(x)" is not a function of bounded variation on "[a,b]". The received contradiction proves the theorem.


Theorem 2. If "c\in(a,b)" then "V_a^b(f)=V_a^c(f)+V_c^b(f)".

Proof. Let "P_1=\{x_0, x_1, \dots,x_{n_1}\}\in Part[a,c]", "P_2=\{y_0, y_1, \dots,y_{n_2}\}\in Part[c,b]". Put "P=P_1\cup P_2\in Part[a,b]". Then

"V(f;P)=\sum\limits_{i=1}^{n_1}|f(x_i)-f(x_{i-1})|+\sum\limits_{j=1}^{n_2}|f(y_j)-f(y_{j-1})|=V(f;P_1)+V(f;P_2)"

Hence

(*) "V_a^b(f)=\sup\{V(f;P): P\in Part[a,b]\}\geq"

"\sup\{V(f;P_1)+V(f;P_2) | P_1\in Part[a,c], P_2\in Part[c,b]\}="

"=V_a^c(f)+V_c^b(f)"

By now, let "P\in Part[a,b]" be arbitrary. Put "P'=P\cup\{c\}".

If "c\in P", then "P'=P" and"V(f;P')=V(f;P)".

If "c\notin P" then "c\in(x_k, x_k+1)" for some "k" and, hence, "V(f;P')-V(f;P)=|f(c)-f(x_k)|+|f(x_{k+1})-f(c)|-|f(x_{k+1})-f(x_k)|\geq 0" by the triangle inequality.

Put "LP=P'\cap [a,c]\in Part[a,c]", "RP=P'\cap [c,b]\in Part[c,b]". Then "P'=LP\cup RP" and "LP\cap RP=\{c\}". For such partitions it was proved above that "V(f; P')=V(f; LP)+V(f; RP)"

Therefore

"V(f; P)\leq V(f; P')=V(f; LP)+V(f; RP)\leq V_a^c(f)+V_c^b(f)"

Taking the supremum by "P" , we obtain

"V_a^b(f)\leq V_a^c(f)+V_c^b(f)"

Combining this with (*), we get "V_a^b(f)= V_a^c(f)+V_c^b(f)".

Q.E.D.